# Empowering Scientific Workflows with Federated Agents

### v2

@@ 6. Related Work
A workflow is a structured sequence of tasks, typically a directed acyclic graph (DAG), designed to achieve a specific goal, often involving data transformation, analysis, or computational modeling. Frameworks for building workflows take many forms. Parallel computing libraries, such as Dask (Rocklin, 2015) and Ray (Moritz et al., 2018), provide mechanisms for executing functions in parallel across local resources or distributed systems. Similarly, workflow management systems (WMSs) can execute tasks in parallel but also provide mechanisms for defining, optimizing, and monitoring DAG execution (e.g., Airflow (Apache, 2015), Fireworks (Jain et al., 2015), Makeflow (Albrecht et al., 2012), Nextflow (Di Tommaso et al., 2017), Parsl (Babuji et al., 2019), Pegasus (Deelman et al., 2015), Swift (Wilde et al., 2009)). WMSs can be differentiated by how dependency graphs are defined (Pauloski et al., 2024): static configurations files, such as CWL (Crusoe et al., 2022), XML, or YAML; general purpose languages (GPLs); domain specific languages (DSLs); or procedurally through the dynamic execution of a program. The class of workflows supported by these frameworks have two key limitations that we address: tasks are assumed to be pure (i.e., no side-effects) and programs are static, i.e., they cannot adapt to changing environments over time.
@@
Actors are computational entities that enable concurrent computing through message passing (Hewitt et al., 1973). In response to a message, an actor can alter its local state, send messages to other actors, and create new actors. No global state means locks and other synchronization primitives are not required. Actors can enable stateful computations within traditionally stateless programming models, and are supported in parallel computing frameworks (e.g., Akka (Lightbend, 2009), Dask, Orleans (Bernstein et al., 2014), Ray) and function-as-a-service (FaaS) platforms (e.g., Abaco (Garcia et al., 2020), Azure Service Fabric (Microsoft, 2017), PraaS (Copik et al., 2022)). Actor models have been investigated as alternatives for designing computational workflows where communication and coordination are decoupled (Bowers and Ludäscher, 2005). Our system extends the actor model to support autonomous behaviors and federated deployments.
@@
Multi-agent systems can enhance or automate scientific processes. Early work investigated cooperative agent environments for distributed problem solving with minimal human intervention (Drashansky et al., 1995, 1999). Recent work focuses on improving the reasoning capabilities of LLM-backed agents through ontological knowledge graphs and multi-agent systems (Ghafarollahi and Buehler, 2024) and tool-augmented LLMs (Ma et al., 2024). Increasingly popular is the use of multi-agent conversations, in which multiple role-specialized agents interact to collaborate, coordinate, or compete towards goals (Wu et al., 2023). These systems enhance LLM-based tools through better reasoning (Du et al., 2023), validation (Wu et al., 2024), and divergent thinking (Liang et al., 2024), prompting rapid development of frameworks such as LangGraph (LangChain, 2024), Microsoft AutoGen (Wu et al., 2023), OpenAI Swarm (OpenAI, 2024), and Pydantic Agents (Pydantic, 2024). Subsequently, interest in standardizing agent protocols has developed. Anthropic’s Model Context Protocol (MCP) (Anthropic, 2024) defines structured interaction between humans/tools and AI models. Google’s Agent2Agent (A2A) Protocol (Google, 2025) focuses on structured interaction between autonomous agents; each agent serves an HTTP endpoint which is impracticable for many scientific workflows. Multi-agent conversations can proxy scientists in iterative scientific processes—brainstorming ideas, planning experiments, and reasoning about results (Boiko et al., 2023; Wang et al., 2023; Gao et al., 2024; Google Research, 2025)—but these aforementioned systems are designed for local or cloud-native applications and lack the features necessary to deploy agents across federated research infrastructure. We focus on the systems-level challenges of representing and deploying diverse agent types and agentic workflows across heterogeneous environments rather than the applied use of LLMs for workflow steering.
7. Conclusion & Future Work
@@
Advancements in AI, coupled with concurrent advancements in self-driving laboratories, high performance computing, and research data management, open the door for truly autonomous scientific discovery. Realizing this grand vision requires mechanisms for the seamless and dynamic integration of research software and infrastructure. To that end, we introduced Academy, a middleware for developing agentic workflows that engage multi-agent systems spanning federated research infrastructure. This framework enables scalable and flexible orchestration of intelligent agents across heterogeneous resources. We presented solutions to three key challenges: representing and programming agents; communicating among agents; and executing agents across diverse resources. Our evaluations demonstrate that Academy can support high-performance workflows, and three case studies highlight the advantages of agentic workflow design.
In future work, we will explore scoped authentication to control which agents can invoke others, enabling the creation of agent marketplaces where access can be granted, revoked, or delegated. We also plan to expand agent discovery with additional metadata to support AI-steered workflows in which LLMs autonomously identify and use available agents. Recording the relative ordering of agent events (i.e., messages received and state transitions), as in Instant Replay (Leblanc and Mellor-Crummey, 1987), can support provenance within agentic workflows. Last but not least, we will work across scientific research communities to assemble agents for different purposes, and with research facilities to identify obstacles to agent use that may motivate further developments in Academy.
8. Acknowledgments

### v3

@@ VII Related Work
A workflow is a structured sequence of tasks, typically a directed acyclic graph (DAG), designed to achieve a specific goal, often involving data transformation, analysis, or computational modeling. Frameworks for building workflows take many forms. Parallel computing libraries, such as Dask [53] and Ray [54], provide mechanisms for executing functions in parallel across local resources or distributed systems. Similarly, workflow management systems (WMSs) can execute tasks in parallel but also provide mechanisms for defining, optimizing, and monitoring DAG execution (e.g., Airflow [55], Fireworks [56], Makeflow [57], Nextflow [58], Parsl [32], Pegasus [59], Swift [60]). WMSs can be differentiated by whether dependency graphs are defined [61] with static configurations files, such as CWL [62], XML, or YAML; general purpose languages (GPLs); domain specific languages (DSLs); or procedurally through the dynamic execution of a program. The class of workflows supported by these frameworks have two key limitations that we address: tasks are assumed to be pure (i.e., no side-effects) and programs are static, i.e., they cannot adapt to changing environments over time.
@@
Actors are computational entities that enable concurrent computing through message passing [20]. In response to a message, an actor can alter its local state, send messages to other actors, and create new actors. No global state means locks and synchronization primitives are not required. Actors can enable stateful computations within traditionally stateless programming models, and are supported in parallel computing frameworks (e.g., Akka [63], Dask, Orleans [64], Ray) and function-as-a-service (FaaS) platforms (e.g., Abaco [65], Azure Service Fabric [66], PraaS [67]). Actor models have been investigated as alternatives for designing computational workflows where communication and coordination are decoupled [68]. Our system extends the actor model to support autonomous behaviors and federated deployments.
@@
Multi-agent systems can enhance or automate scientific processes. Early work investigated cooperative agent environments for distributed problem solving with minimal human intervention [69, 70]. Recent work focuses on improving the reasoning capabilities of LLM-backed agents through ontological knowledge graphs and multi-agent systems [71] and tool-augmented LLMs [72]. Increasingly popular is the use of multi-agent conversations, in which multiple role-specialized agents interact to collaborate, coordinate, or compete towards goals [11]. These systems enhance LLM-based tools through better reasoning [73], validation [74], and divergent thinking [75], prompting rapid development of frameworks such as LangGraph [12], Microsoft AutoGen [11], OpenAI Swarm [13], and Pydantic Agents [76]. Subsequently, interest in standardizing agent protocols has developed. Anthropic’s Model Context Protocol (MCP) [77] defines structured interaction between humans/tools and AI models. Google’s Agent2Agent (A2A) Protocol [78] focuses on structured interaction between autonomous agents; each agent serves an HTTP endpoint which is impracticable for many scientific workflows. Multi-agent conversations can proxy scientists in iterative scientific processes—brainstorming ideas, planning experiments, and reasoning about results [79, 9, 10, 80]—but these aforementioned systems are designed for local or cloud-native applications and lack the features necessary to deploy agents across federated research infrastructure. We focus on the systems-level challenges of representing and deploying diverse agent types and agentic workflows across heterogeneous environments rather than the applied use of LLMs for workflow steering.
VIII Conclusion & Future Work
@@
Advancements in AI, coupled with concurrent advancements in self-driving laboratories, high performance computing, and research data management, open the door for truly autonomous scientific discovery. Realizing this grand vision requires mechanisms for the seamless and dynamic integration of research software and infrastructure. To that end, we introduced Academy, a middleware for developing agentic workflows that engage multi-agent systems spanning federated research infrastructure. This framework enables scalable and flexible orchestration of intelligent agents across heterogeneous resources. We presented solutions to three key challenges: representing and programming agents; communicating among agents; and executing agents across diverse resources. Our evaluations demonstrate that Academy can support high-performance workflows, and four case studies highlight the advantages of agentic workflow design.
IX Acknowledgments